### v1

# Demystifying Spectral Feature Learning for Instrumental Variable Regression

Abstract
We address the problem of causal effect estimation in the presence of hidden confounders, using nonparametric instrumental variable (IV) regression. A leading strategy employs spectral features - that is, learned features spanning the top eigensubspaces of the operator linking treatments to instruments. We derive a generalization error bound for a two-stage least squares estimator based on spectral features, and gain insights into the method’s performance and failure modes. We show that performance depends on two key factors, leading to a clear taxonomy of outcomes. In a good scenario, the approach is optimal. This occurs with strong spectral alignment, meaning the structural function is well-represented by the top eigenfunctions of the conditional operator, coupled with this operator’s slow eigenvalue decay, indicating a strong instrument. Performance degrades in a bad scenario: spectral alignment remains strong, but rapid eigenvalue decay (indicating a weaker instrument) demands significantly more samples for effective feature learning. Finally, in the ugly scenario, weak spectral alignment causes the method to fail, regardless of the eigenvalues’ characteristics. Our synthetic experiments empirically validate this taxonomy.
1 Introduction
We study the nonparametric instrumental variable (NPIV) model [1, 30]
[…]
where $X$ is the treatment, $Y$ the outcome, $Z$ the instrument, $U$ an unobserved confounder, and $h_{0}\in L_{2}(X)$ is the structural function to be learned, where we denoted $L_{2}(X)$ the $L_{2}$ space corresponding to the distribution $\mathbb{P}_{X}$. We assume no common confounder of $Z$ and $Y$. To illustrate this setting, consider a simplified example inspired by [19]. Suppose we want to determine how ticket price ($X$) affects the number of flight tickets sold ($Y$). A potential confounder ($U$) could be a major event (e.g., a large conference or festival) occurring at the destination. This event would likely increase demand for flights, and airline pricing algorithms would result in more expensive seats being sold; the confounder $U$ influences both $X$ and $Y$. A naive regression of sales on price might then misleadingly suggest that higher prices lead to higher sales. This occurs because the analysis fails to account for the event, which independently drives up both demand (sales) and prices. To obtain a less biased estimate of the price effect, we could use an instrument ($Z$), such as the cost of jet fuel. The cost of fuel is a plausible instrument because: (i) it likely affects ticket prices ($X$) as airlines adjust fares based on operational costs, and (ii) it is unlikely to be directly correlated with whether a major event ($U$) is happening at the destination, thus satisfying $\mathbb{E}\left[U\mathrel{}\middle|\mathrel{}Z\right]=0$. By using fuel cost as an instrument, we can isolate the variation in ticket prices that is independent of the unobserved demand shock caused by the event, allowing for a more accurate estimation of the structural function $h_{0}$.
Central to the NPIV model is the conditional expectation operator $\mathcal{T}:L_{2}(X)\rightarrow L_{2}(Z)$, defined by $\mathcal{T}h(Z)=\mathbb{E}[h(X)\mid Z]$. Estimating $h_{0}$ amounts to inverting this operator from the following integral equation
$\mathcal{T}h_{0}=r_{0},\quad\text{where }r_{0}(Z)\doteq\mathbb{E}[Y\mid Z].$ (2)
In practice, the operator $\mathcal{T}$ is unknown and must be estimated from data. Moreover, even if $\mathcal{T}$ were known, its inverse is typically unbounded, making recovery of $h_{0}$ an ill-posed inverse problem that requires regularization.
A classical approach to solving NPIV is nonlinear two-stage least squares (2SLS), which projects $X$ and $Z$ into suitable feature spaces and performs regression in these representations. Let $\varphi$ and $\psi$ denote feature maps for $X$ and $Z$, respectively. In Stage 1, one estimates the conditional expectation $\mathbb{E}[\varphi(X)\mid Z]$ by regressing $\varphi(X)$ on $\psi(Z)$, yielding an approximation of the form $\mathbb{E}[\varphi(X)\mid Z]\approx A\psi(Z)$, for some linear map $A$. This step isolates the component of $\varphi(X)$ that is predictable from $Z$, filtering out variation due to the unobserved confounder. In Stage 2, the outcome $Y$ is regressed on the predicted features $\mathbb{E}[\varphi(X)\mid Z]$, yielding an estimator of the structural function $h_{0}$, which is modeled as a linear combination $h_{0}(X)\approx\beta^{\mathsf{\scriptscriptstyle T}}\varphi(X)$, for some coefficient vector $\beta$. Given that the feature maps $\varphi$ and $\psi$ can be fixed (e.g., polynomials, splines, or radial basis functions from a kernel) or learned from data (e.g., using neural networks), a natural question arises: What specific information must these features encode to enable efficient estimation?
The algorithm recently proposed by [34], which learns features by minimizing a spectral contrastive loss, suggests that the answer lies in the spectral structure of $\mathcal{T}$. In this paper, we rigorously investigate this perspective. We show that the performance of this approach hinges on two key factors: the spectral alignment of the structural function $h_{0}$ with the operator $\mathcal{T}$, and the rate of its singular value decay. Namely, we identify three distinct regimes governing the effectiveness of spectral methods for causal inference: [^1] The “good–bad–ugly” terminology is borrowed from [23], who used it to describe performance regimes in Koopman operator learning. We adopt the same naming convention in a different context: spectral feature learning for causal inference. The phrase itself, of course, originates from the classic [26].
  1. 1.
    The good regime, where the structural function aligns strongly with the leading eigenspaces of the conditional expectation operator $\mathcal{T}$, and the singular values of $\mathcal{T}$ decay slowly;
  2. 2.
    The bad regime, where spectral alignment holds, but rapid singular value decay leads to instability and degraded estimation quality;
  3. 3.
    The ugly regime, where the structural function is misaligned with the top eigenspaces of $\mathcal{T}$, rendering spectral methods ineffective regardless of decay rate.
Overview of contributions. We investigate the effectiveness of spectral features in NPIV estimation, grounding their empirical success in a principled theoretical analysis. Specifically:
$\bullet$ By leveraging an upper bound on the generalization error for the 2SLS estimator and specializing it to spectral features, we characterize the conditions under which each regime arises.
$\bullet$ We clarify the theoretical foundation of the state-of-the-art causal estimation algorithm of [34] by showing that it implements a 2SLS estimator with learned features that explicitly target the leading eigenspaces of the conditional expectation operator $\mathcal{T}$. Although [34] suggested a heuristic link to the Singular Value Decomposition (SVD) of $\mathcal{T}$, we make this connection explicit by proving that their contrastive objective corresponds to a Hilbert-Schmidt approximation of $\mathcal{T}$.
$\bullet$ We empirically validate the “good–bad–ugly” taxonomy through synthetic experiments designed to isolate the effects of spectral alignment and singular value decay on generalization error.
Related work. Instrumental variable regression is often implemented as a two-stage procedure, where the key design choice is the feature representation. Fixed features include polynomials, splines, or kernel methods [3, 6, 7, 29, 33], while learned features are either derived from the conditional expectation operator $\mathcal{T}$ [34] or jointly optimized with the 2SLS objective [22, 31, 37]. Some approaches replace first-stage regression with conditional density estimation [10, 17, 19, 27, 32].
An alternative line of work reformulates IV regression as a saddle-point optimization problem to avoid the bias introduced by nested conditional expectation estimation [2, 12, 28, 36, 39]. These methods frame estimation as a min-max game between a candidate solution $h$ and a dual witness function enforcing the IV moment condition. This approach sidesteps explicit conditional expectation estimation but relies on selecting tractable, fixed function classes to ensure the well-posedness of the optimization problem.
Features can be learned in various ways. One line of work learns them jointly with the regression objective, as in [15, 22, 31, 37], who propose end-to-end frameworks that simultaneously optimize over the feature representations and second-stage parameters. In contrast, [34, 36] propose to learn features that reflect the eigenstructure of $\mathcal{T}$, independently of the outcome $Y$, using a spectral contrastive loss. This may appear suboptimal, since incorporating $Y$ could, in principle, yield more informative representations, however it avoids the optimization difficulties inherent to nonconvex, joint objectives involving all three variables $(X,Y,Z)$. This decoupled spectral approach was shown in certain experimental settings to outperform end-to-end alternatives, as demonstrated in [34]. Our work places this empirical finding in context, since it likely arises due to the experiments in that work satisfying the “good” regime in the present work. We provide further discussion in Appendix B.
Structure of the paper. Section 2 introduces the notation and preliminaries. Section 3 reviews the Sieve 2SLS estimator and presents generalization bounds highlighting the role of spectral features. Section 4 shows how to learn such features via a contrastive loss and connects this to the method of [34]. Section 5 provides synthetic experiments that validate our theoretical taxonomy. Proofs are deferred to the Appendix.
2 Preliminaries
2.1 Background
[…]
Operators on Hilbert Spaces. Let $H$ be a Hilbert space. For a bounded linear operator $A$ acting on $H$, we denote its operator norm by $\|A\|_{\mathrm{op}}$, its Hilbert–Schmidt norm by $\left\|A\right\|_{\scriptscriptstyle\mathrm{HS}}$, its Moore–Penrose inverse by $A^{\dagger}$, and its adjoint by $A^{\star}$. For finite-dimensional operators, the Hilbert–Schmidt norm coincides with the Frobenius norm. We denote the range and null space of $A$ by $\mathcal{R}(A)$ and $\mathcal{N}(A)$, respectively. Given a closed subspace $M\subseteq H$, we write $M^{\perp}$ for its orthogonal complement, $\overline{M}$ for its closure, and $\Pi_{M}$ for the orthogonal projection onto $M$. The orthogonal projection onto $M^{\perp}$ is denoted by $(\Pi_{M})_{\perp}\doteq I_{H}-\Pi_{M}=\Pi_{M^{\perp}}$. For functions $f,h\in L_{2}(X)$, $g\in L_{2}(Z)$, the rank-one operator $g\otimes f$ is defined as $(g\otimes f)(h)=\langle h,f\rangle_{L_{2}(X)}g$. This generalizes the standard outer product. For vectors $x\in\mathbb{R}^{d}$, we write $\|x\|_{\ell_{2}}$ for the Euclidean norm.
Data Splitting and Empirical Expectations. Given $n,m\geq 0$, we consider two independent datasets: an unlabeled dataset, $\tilde{\mathcal{D}}_{m}=\{(\tilde{z}_{i},\tilde{x}_{i})\}_{i=1}^{m}$, used to learn features for $X$ and $Z$, and a labeled dataset, $\mathcal{D}_{n}=\{(z_{i},x_{i},y_{i})\}_{i=1}^{n}$, used to estimate the structural function. For $R\in\{X,Z,Y,(X,Z),(X,Y),(Z,Y),(X,Z,Y)\}$ and a measurable function $f$, we define the empirical expectation over $\tilde{\mathcal{D}}_{m}$ as $\hat{\mathbb{E}}_{m}[f(R)]\doteq\frac{1}{m}\sum_{i=1}^{m}f(\tilde{r}_{i})$, and similarly for $\hat{\mathbb{E}}_{n}$ on $\mathcal{D}_{n}$.
Feature Maps and Projection Operators. Let $d\geq 1$, and let $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ be a feature map with components $\varphi_{i}\in L_{2}(X)$ for $i=1,\dots,d$. We define $\mathcal{H}_{\varphi,d}\doteq\operatorname{span}\{\varphi_{1},\ldots,\varphi_{d}\}$, and let $\Pi_{\varphi,d}$ be the $L_{2}-$orthogonal projection onto this space. Analogous definitions apply for feature maps $\psi:\mathcal{Z}\to\mathbb{R}^{d}$, yielding $\mathcal{H}_{\psi,d}$ and $\Pi_{\psi,d}$.
2.2 On the Conditional Expectation Operator: NPIV as an Ill-posed Inverse Problem
We begin by defining the conditional expectation operator. Let
$\mathcal{T}:L_{2}(X)\to L_{2}(Z),\quad h\mapsto\mathbb{E}[h(X)\mid Z]\,.$
As highlighted by Eq. 2, this operator plays a central role in the NPIV model. It encodes the conditional distribution of $X$ given $Z$ and provides a convenient lens through which to analyze how efficiently one can estimate $h_{0}$. The NPIV equation admits a solution if and only if $r_{0}\in\mathcal{R}(\mathcal{T})$, and is identifiable if and only if $\mathcal{T}$ is injective. We therefore adopt the following assumption.
Assumption 1.
$\mathcal{T}$ is injective and there exists a solution to the NPIV problem, i.e., $\mathcal{T}^{-1}(\{r_{0}\})\neq\emptyset$.
The existence part is essential: without it, there is no structural function $h_{0}$ satisfying the model in Eq. 1. While injectivity can in principle be relaxed, this comes at the cost of recovering a different object. Specifically, if $h_{0}$ is not identifiable, one may instead recover the minimum-norm solution to Eq. 2, namely $h_{*}\doteq\mathcal{T}^{\dagger}r_{0}$ [14]. [^2] This corresponds to the unique element in the pre-image of $r_{0}$ under $\mathcal{T}$ that minimizes the $L_{2}$-norm and is equivalently characterized as the unique element of the set $\mathcal{T}^{-1}(\{r_{0}\})\cap\mathcal{N}(\mathcal{T})^{\perp}$ [2]. We assume injectivity henceforth to simplify the discussion.
Eq. 2 is an inverse problem in which both the operator $\mathcal{T}$ and the right-hand side $r_{0}$ are unknown. This inverse problem is typically ill-posed, as $\mathcal{T}^{-1}$ is generally not continuous. [^3] In particular, $\mathcal{T}^{-1}$ is discontinuous if $\mathcal{T}$ is compact and not finite rank. Hence, even if $\mathcal{T}$ were known, an approximation $\hat{r}\approx r_{0}$ would not ensure that $\mathcal{T}^{-1}\hat{r}$ is close to $h_{0}=\mathcal{T}^{-1}r_{0}$. In such settings, regularization is required to obtain stable approximate inverses. Two common regularization approaches are Tikhonov regularization, where $\mathcal{T}^{-1}\approx(\mathcal{T}^{\star}\mathcal{T}+\lambda I)^{-1}\mathcal{T}^{\star}$, $\lambda>0$, and spectral cut-off, which truncates the Singular Value Decomposition (SVD) [13].
When $\mathcal{T}$ is compact (and therefore it admits a countable sequence of singular values), the degree of ill-posedness of the inverse problem is characterized by the decay rate of these singular values. Since the singular values of $\mathcal{T}^{-1}$ are the reciprocals of those of $\mathcal{T}$, faster decay implies worse ill-posedness. The NPIV problem can be viewed as particularly challenging, as in addition to the ill-posedness of the operator $\mathcal{T}$, one must also estimate it from data. We now introduce a sufficient condition for compactness of $\mathcal{T}$.
Assumption 2.
The joint distribution $\pi_{XZ}$ is dominated by the product measure $\pi_{X}\otimes\pi_{Z}$, and its Radon–Nikodym derivative, or “density”, $p\doteq\frac{\mathop{}\!\mathrm{d}\pi_{XZ}}{\mathop{}\!\mathrm{d}\pi_{X}\otimes\pi_{Z}}$ belongs to $L_{2}(\mathcal{X}\times\mathcal{Z},\pi_{X}\otimes\pi_{Z})$.
Assumption 2 is standard in the NPIV literature (Assumption A.1 in [10]). It may fail when $X$ and $Z$ share variables or are deterministically related. For example, if $Z=(W,V)$ and $X=V$, then the joint distribution is not dominated by the product measure.
Proposition 1.
Under Assumption 2, $\mathcal{T}$ is a Hilbert–Schmidt operator and thus compact.
This result is well known in the inverse problems literature; we provide a proof in Appendix A.1.
Compact Decomposition and Eigensubspaces. As $\mathcal{T}$ is a conditional expectation operator, it is non-expansive, i.e., $\|\mathcal{T}\|_{\text{op}}\leq 1$. Under Assumption 2, $\mathcal{T}$ is compact, and therefore admits a SVD
$\mathcal{T}=\sum_{i\geq 1}\sigma_{i}u_{i}\otimes v_{i}\,,$ (3)
where $\{u_{i}\}\subset L_{2}(Z)$ and $\{v_{i}\}\subset L_{2}(X)$ are orthonormal systems (left and right singular functions, respectively) spanning $\mathcal{N}(\mathcal{T})^{\perp}$ and $\overline{\mathcal{R}(\mathcal{T})}$ respectively, and $\sigma_{i}$ are the nonnegative singular values in nonincreasing order. The leading singular triplet is known explicitly: $\sigma_{1}=1$, $u_{1}=\mathbbm{1}_{\mathcal{Z}}$ and $v_{1}=\mathbbm{1}_{\mathcal{X}}$ corresponding to the constant functions. If $\mathcal{T}$ is infinite-rank, then $\sigma_{i}\to 0$ as $i\to\infty$. Under the assumption that $\mathcal{T}$ is injective (Assumption 1), we have $\mathcal{N}(\mathcal{T})=\{0\}$, hence $\{v_{i}\}_{i\geq 1}$ forms an orthonormal basis (ONB) of $L_{2}(X)$ and $\sigma_{i}>0$ for all $i\geq 1$. Finally, since $\mathcal{T}$ is Hilbert–Schmidt, $\sum_{i\geq 1}\sigma_{i}^{2}<\infty$. For any $d\geq 1$, define the leading eigensubspaces of $\mathcal{T}$ as
$\mathcal{U}_{d}$ $\doteq\text{span}\{u_{1},\ldots,u_{d}\},\quad\Pi_{\mathcal{Z},d}:\text{ orthogonal projection onto }\mathcal{U}_{d},$
$\mathcal{V}_{d}$ $\doteq\text{span}\{v_{1},\ldots,v_{d}\},\quad\Pi_{\mathcal{X},d}:\text{ orthogonal projection onto }\mathcal{V}_{d}.$
In addition, define $\mathcal{T}_{d}\doteq\sum_{i=1}^{d}\sigma_{i}u_{i}\otimes v_{i}$ the rank-$d$ truncation of $\mathcal{T}$. To avoid ambiguity in the definition of $\mathcal{U}_{d},\mathcal{V}_{d}$ and $\mathcal{T}_{d}$, we assume that $\sigma_{d}>\sigma_{d+1}$.
3 Solving NPIV with 2SLS
To understand how and when spectral features should be used, we first recall results from the Sieve 2SLS literature. This gives us a clear picture of what terms are to be controlled in order to achieve good generalization, which we will use to demonstrate the effectiveness of spectral features and reveal our good-bad-ugly taxonomy.
3.1 Sieve 2SLS estimator
We start with the following characterization of the structural function:
$h_{0}=\operatorname*{arg\,min}_{h\in L_{2}\left(X\right)}\,\mathbb{E}\left[\left(Y-\mathcal{T}h\left(Z\right)\right)^{2}\right]\,,$
where the minimizer, $h_{0}$, is unique under Assumption 1. A popular strategy to estimate $h_{0}$ is to consider a hypothesis space $\mathcal{H}_{\mathcal{X}}\subset L_{2}\left(X\right)$ and a procedure to estimate the action of $\mathcal{T}$: $\hat{\mathcal{T}}h\approx\mathcal{T}h$ for all $h\in\mathcal{H}_{\mathcal{X}}$, in order to obtain
$\hat{h}\in\operatorname*{arg\,min}_{h\in\mathcal{H}_{\mathcal{X}}}\,\hat{\mathbb{E}}_{n}\left[\left(Y-\hat{\mathcal{T}}h\left(Z\right)\right)^{2}\right]\,,$ (4)
possibly subject to regularization. In the algorithm 2SLS, we consider features on both $\mathcal{X}$ and $\mathcal{Z}$, and both steps (estimating $h\mapsto\mathcal{T}h$ and solving Eq. 4) are carried out via linear regression in feature space. The two sets of features may be infinite-dimensional—e.g., using Reproducing Kernel Hilbert Spaces (RKHS)—or finite-dimensional—e.g., using splines, wavelets, or neural network features. When the feature dimension grows with sample size, we obtain a sieve. Since our focus is on neural network-based features, we present the finite-dimensional version, noting that the ideas naturally extend to infinite-dimensional cases with proper regularization [33, 29]. We refer to Appendix B for a description of other 2SLS methods.
Consider a feature map $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ such that $\mathcal{H}_{\mathcal{X}}\doteq\{x\mapsto\theta^{\mathsf{\scriptscriptstyle T}}\varphi\left(x\right),\theta\in\mathbb{R}^{d}\}$ is included in $L_{2}\left(X\right)$. Then for any $h=\theta^{\mathsf{\scriptscriptstyle T}}\varphi\left(\cdot\right)\in\mathcal{H}_{\mathcal{X}}$, $\mathcal{T}h\left(Z\right)=\theta^{\mathsf{\scriptscriptstyle T}}\mathbb{E}\left[\varphi\left(X\right)\mid Z\right]\,.$ Therefore, estimating $F_{\star}(Z)\doteq\mathbb{E}[\varphi(X)\mid Z]$ allows us to estimate $\mathcal{T}h$ for any $h\in\mathcal{H}_{\mathcal{X}}$. Let $\psi:\mathcal{Z}\to\mathbb{R}^{d}$ be another feature map with $\mathcal{H}_{\mathcal{Z}}\doteq\{z\mapsto\alpha^{\mathsf{\scriptscriptstyle T}}\psi(z),\alpha\in\mathbb{R}^{d}\}\subset L_{2}(Z)$. We estimate $F_{\star}$ with vector-valued linear regression in feature space by approximating $F_{\star}$ with $A\psi(\cdot)$, where $A\in\mathbb{R}^{d\times d}$, solving
$\hat{A}\doteq\hat{\mathbb{E}}_{n}[\varphi(X)\psi(Z)^{\mathsf{\scriptscriptstyle T}}]\hat{\mathbb{E}}_{n}[\psi(Z)\psi(Z)^{\mathsf{\scriptscriptstyle T}}]^{\dagger}\in\operatorname*{arg\,min}_{A\in\mathbb{R}^{d\times d}}\hat{\mathbb{E}}_{n}[\|\varphi(X)-A\psi(Z)\|_{\ell_{2}}^{2}]\,.$
The pseudo-inverse yields the minimum-norm solution in the case of non-uniqueness. Regularized versions, such as Tikhonov-regularized least squares (also known as ridge regression), can be obtained by penalizing with $\eta\left\|A\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}$ for $\eta>0$. Computing $\hat{A}$ is referred to as the first stage of the 2SLS algorithm. For the second stage, for any $h=\theta^{\mathsf{\scriptscriptstyle T}}\varphi(\cdot)\in\mathcal{H}_{\mathcal{X}}$, we define $\hat{\mathcal{T}}h(Z)\doteq\theta^{\mathsf{\scriptscriptstyle T}}\hat{F}(Z)$ with $\hat{F}(Z)=\hat{A}\psi(Z)$. Plugging this into Eq. 4 yields $\hat{h}=\hat{\theta}^{\mathsf{\scriptscriptstyle T}}\varphi(\cdot)$ with
$\hat{\theta}$ $\doteq\hat{\mathbb{E}}_{n}[\hat{F}(Z)\hat{F}(Z)^{\mathsf{\scriptscriptstyle T}}]^{\dagger}\hat{\mathbb{E}}_{n}[\hat{F}(Z)Y]\in\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\hat{\mathbb{E}}_{n}[(Y-\theta^{\mathsf{\scriptscriptstyle T}}\hat{F}(Z))^{2}]\,.$ (5)
Similarly to the first stage, we can also consider a regularized version of the second stage, where we penalize with $\lambda\|\theta\|_{\ell_{2}}^{2}$ for $\lambda>0$. $\hat{h}$ is the sieve NPIV estimator [3, 7, 6].
3.1.0.0.1 2SLS versus saddle-point estimator.
Based on features $\varphi$ and $\psi$, [39, 34] proposed an estimator in a saddle-point form: $\hat{h}_{bis}=\hat{\theta}_{bis}^{\mathsf{\scriptscriptstyle T}}\varphi(\cdot)$, with
$\hat{\theta}_{bis}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\max_{\alpha\in\mathbb{R}^{d}}\hat{\mathbb{E}}_{n}[\alpha^{\mathsf{\scriptscriptstyle T}}\psi(Z)\left(Y-\theta^{\mathsf{\scriptscriptstyle T}}\varphi(X)\right)-0.5\cdot(\alpha^{\mathsf{\scriptscriptstyle T}}\psi(Z))^{2}]+\lambda\|\theta\|_{\ell_{2}}^{2}\,.$
Albeit looking different, solving this min-max problem in closed-form leads to $\hat{\theta}_{bis}=\hat{\theta}$ (when we introduce regularization in Eq. 5), and thus $\hat{h}_{bis}=\hat{h}$. The saddle-point form could be useful for deriving convergence rates, see [2] for example, as it allows us to use the theory of saddle-point problems. However, it is not necessary for the implementation of the algorithm.
3.2 Convergence rates for Sieve 2SLS
In order to present convergence rates for Sieve 2SLS, we introduce the sieve measure of ill-posedness.
Definition 1.
Given $d\geq 1$ and $\mathcal{H}_{\varphi,d}=\operatorname{span}\{\varphi_{1},\ldots,\varphi_{d}\}$, the sieve measure of ill-posedness is defined as
$\tau_{\varphi,d}\doteq\sup_{h\in\mathcal{H}_{\varphi,d},~{}h\neq 0}\frac{\|h\|_{L_{2}(X)}}{\|\mathcal{T}h\|_{L_{2}(Z)}}=\left(\inf_{h\in\mathcal{H}_{\varphi,d},~{}\|h\|_{L_{2}}=1}\|\mathcal{T}h\|_{L^{2}(Z)}\right)^{-1}.$
This quantity can be interpreted as the operator norm of $\mathcal{T}^{-1}$ restricted to $\mathcal{T}(\mathcal{H}_{\varphi,d})$; that is, $\tau_{\varphi,d}=\sup_{g\in\mathcal{T}(\mathcal{H}_{\varphi,d}),~{}h\neq 0}\|\mathcal{T}^{-1}g\|_{L_{2}(X)}/\|g\|_{L_{2}(Z)}.$ While the full operator $\mathcal{T}^{-1}$ is typically unbounded, the restricted norm $\tau_{\varphi,d}$ remains finite. It captures the degree of ill-posedness specific to the chosen hypothesis space $\mathcal{H}_{\varphi,d}$. As $d$ increases and the hypothesis class becomes richer, $\tau_{\varphi,d}$ increases. Equipped with this quantity, we now introduce stability conditions that will allow us to establish generalization bounds for Sieve 2SLS.
Assumption 3.
(i) $\sup_{h\in\mathcal{H}_{\varphi,d},~{}\|h\|_{L_{2}}=1}\left\|\left(\Pi_{\psi,d}\right)_{\perp}\mathcal{T}h\right\|_{L^{2}(Z)}=o_{d}\left(\tau_{\varphi,d}^{-1}\right)$; (ii) there is a constant $C>0$ such that $\left\|\mathcal{T}\left(\Pi_{\varphi,d}\right)_{\perp}h_{0}\right\|_{L^{2}(Z)}\leq C\times\tau_{\varphi,d}^{-1}\times\left\|\left(\Pi_{\varphi,d}\right)_{\perp}h_{0}\right\|_{L^{2}(X)}$.
Assumption 3-(i) is a condition on the approximation properties of the features used for the instrument space. It says the image of $\mathcal{H}_{\varphi,d}$ through $\mathcal{T}$ lies almost entirely inside $\mathcal{H}_{\psi,d}$, i.e., the projection onto the orthogonal complement vanishes faster than the inverse ill-posedness. Assumption 3-(ii) is a stability condition sometimes referred as link condition that is standard in the NPIV literature (see Assumption 6 in [3], Assumption 5.2(ii) in [7] and Assumption 3 in [22]). It says that $\mathcal{T}$ becomes increasingly contractive on the component of $h_{0}$ lying outside the sieve space as the sieve dimension increases. The next assumption ensures the conditional variance of the noise is uniformly bounded.
Assumption 4.
There exists $\bar{\sigma}>0$ such that $\mathbb{E}\left[U^{2}\mid Z\right]\leq\bar{\sigma}^{2}<\infty$ almost everywhere.
The following result is an upper bound on the generalization error from [6].
Theorem 1 (Theorem B.1. [6]).
Let $\hat{h}$ be the 2SLS estimator from Eq. 5, using features $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and $\psi:\mathcal{Z}\to\mathbb{R}^{d}$. Suppose Assumptions 1, 3 and 4 hold and $\tau_{\varphi,d}\zeta_{\varphi,\psi,d}\sqrt{(\log d)/n}=o(1)$, where $\zeta_{\varphi,\psi,d}\doteq\max\{\sup_{x}\left\|\mathbb{E}[\varphi(X)\varphi(X)^{\mathsf{\scriptscriptstyle T}}]^{-1/2}\varphi(x)\right\|_{\ell_{2}},\ \sup_{z}\left\|\mathbb{E}[\psi(Z)\psi(Z)^{\mathsf{\scriptscriptstyle T}}]^{-1/2}\psi(z)\right\|_{\ell_{2}}\}$. Then:
$\left\|\hat{h}-h_{0}\right\|_{L^{2}(X)}=O_{p}\left(\left\|(\Pi_{\varphi,d})_{\perp}h_{0}\right\|_{L^{2}(X)}+\tau_{\varphi,d}\sqrt{d/n}\right).$
The first term captures the approximation error, while the second term corresponds to the estimation error arising from finite samples. When standard bases such as cosine, spline, or wavelet functions are used, the approximation error is well-understood [3, 6]. If the sieve ill-posedness $\tau_{\varphi,d}$ grows at a polynomial or exponential rate in $d$, one can choose $d=d(n)$ to balance the two terms optimally. Under such growth conditions and smoothness assumptions on $h_{0}$ (e.g., $h_{0}$ belongs to a Sobolev ball), the resulting convergence rate is minimax optimal.
3.3 Convergence rates for Sieve 2SLS with spectral features
A drawback of using non-adaptive features (such as cosine or spline bases) is that verifying Assumption 3 or characterizing the growth of $\tau_{\varphi,d}$ is generally difficult. In contrast, we now consider features constructed from the SVD of $\mathcal{T}$, for which both conditions can be exactly characterized. Moreover, while estimators based on universal bases can achieve minimax optimality under smoothness assumptions on $\mathcal{T}$ and $h_{0}$, they are often outperformed in practice by estimators using data-dependent features. This highlights that minimax rates, though important, do not capture the full picture of empirical performance. Recall from Eq. 3 the singular value decomposition of $\mathcal{T}$, and the definition of the top-$d$ right eigenspace $\mathcal{V}_{d}$. The following result shows that this subspace achieves the minimal possible measure of ill-posedness among all $d$-dimensional subspaces of $L_{2}(X)$.
Proposition 2 (Lemma 1 [3]).
Let Assumption 2 hold. For any $d\geq 1$, the smallest possible value of $\tau_{\varphi,d}$ is $\tau_{\varphi,d}=\sigma_{d}^{-1}$, achieved when $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$.
Definition 2 (Spectral features).
We say that the features $\varphi$ and $\psi$ are spectral features if
$\mathcal{H}_{\varphi,d}=\operatorname{span}\{v_{1},\ldots,v_{d}\}=\mathcal{V}_{d}\quad\text{and}\quad\mathcal{H}_{\psi,d}=\operatorname{span}\{u_{1},\ldots,u_{d}\}=\mathcal{U}_{d}.$
By Proposition 2, spectral features minimize the sieve measure of ill-posedness, and therefore the estimation error term. This leads to the following corollary.
Corollary 1 (Sieve 2SLS with spectral features).
Let $\hat{h}$ be the 2SLS estimator from Eq. 5 using spectral features. Let Assumptions 1, 2 and 4 hold and $\sigma_{d}^{-1}\zeta_{\varphi,\psi,d}\sqrt{(\log d)/n}=o(1)$. Then:
$\left\|\hat{h}-h_{0}\right\|_{L^{2}(X)}=O_{p}\left(\left\|(\Pi_{\mathcal{X},d})_{\perp}h_{0}\right\|_{L^{2}(X)}+\sqrt{\frac{d}{n\sigma_{d}^{2}}}\right).$
We postpone the proof to Appendix A.2.
The good, the bad, and the ugly. The bound in Corollary 1 decomposes the generalization error into two terms: spectral alignment, encoded in the approximation error $\left\|(\Pi_{\mathcal{X},d})_{\perp}h_{0}\right\|$, which measures how well $h_{0}$ lies in the top-$d$ singular space; and ill-posedness, captured by the inverse singular value $\sigma_{d}^{-1}$. In the good regime, $h_{0}$ lies mostly in $\mathcal{U}_{d}$ and $\sigma_{d}$ decays slowly, so both terms are small. In the bad regime, alignment is favorable but fast spectral decay inflates the estimation error. In the ugly regime, $h_{0}$ is misaligned with the top eigenspaces, so the approximation error dominates and estimation fails regardless of $\sigma_{d}$ or $n$.
4 Sieve 2SLS with learned spectral features
In the previous section, we showed that spectral features minimize the sieve measure of ill-posedness and lead to fast generalization rates when the structural function $h_{0}$ aligns well with the top eigenspaces of the conditional expectation operator $\mathcal{T}$. This motivates learning such features directly from data, especially in the case where the conditional distribution of $X$ given $Z$ is informative (i.e., the instrument is strong) and the alignment with the spectrum of $\mathcal{T}$ is favorable. We now demonstrate that the contrastive learning approach of [34] can be equivalently seen as explicitly targeting the leading eigenspaces of $\mathcal{T}$. Consider the following objective:
$\mathcal{L}_{d}(\varphi,\psi)\doteq\left\|\mathcal{T}_{d}(\varphi,\psi)-\mathcal{T}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2},\qquad\mathcal{T}_{d}(\varphi,\psi)\doteq\sum_{i=1}^{d}\psi_{i}\otimes\varphi_{i},$ (6)
where $\varphi_{i}\in L_{2}(X)$, $\psi_{i}\in L_{2}(Z)$, $i=1,\ldots,d$. As any rank-$d$ operator can be decomposed as $\mathcal{T}_{d}(\varphi,\psi)$, we have the following version of the Eckart–Young–Mirsky Theorem for the Hilbert–Schmidt norm:
Theorem 2 (Eckart–Young–Mirsky Theorem).
Let Assumption 2 hold and let $d\geq 1$ be such that $\sigma_{d}>\sigma_{d+1}$. For all feature maps $\varphi:\mathcal{X}\to\mathbb{R}^{d}$, $\psi:\mathcal{Z}\to\mathbb{R}^{d}$, we have
$\mathcal{L}_{d}(\varphi,\psi)\geq\left\|\mathcal{T}-\mathcal{T}_{d}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}=\sum_{i>d}\sigma_{i}^{2},$
where equality holds if and only if $\mathcal{T}_{d}(\varphi,\psi)=\mathcal{T}_{d}$, in which case $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$, $\mathcal{H}_{\psi,d}=\mathcal{U}_{d}$. $\mathcal{T}_{d}$ is the rank-$d$ truncation of $\mathcal{T}$ introduced in Section 2.
While the Eckart–Young–Mirsky objective in Eq. 6 offers a clean objective to learn spectral features, it appears impractical due to the unknown nature of the operator $\mathcal{T}$. However, this objective admits an equivalent formulation as a spectral contrastive learning loss [35]:
$\mathcal{L}_{d}(\varphi,\psi)=\mathbb{E}_{X}\mathbb{E}_{Z}\left[\left(\varphi(X)^{\mathsf{\scriptscriptstyle T}}\psi(Z)\right)^{2}\right]-2\mathbb{E}_{X,Z}\left[\varphi(X)^{\mathsf{\scriptscriptstyle T}}\psi(Z)\right]+\left\|\mathcal{T}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2},$ (7)
where the last term is independent of $(\varphi,\psi)$. We provide a proof of this equivalence in Appendix A.3. The population loss can be estimated from samples $\tilde{\mathcal{D}}_{m}=\{(\tilde{x}_{i},\tilde{z}_{i})\}_{i=1}^{m}$ as
^$\hat{\mathcal{L}}_{d}(\varphi,\psi)\doteq\frac{1}{m(m-1)}\sum_{i\neq j}\left(\varphi(\tilde{x}_{i})^{\mathsf{\scriptscriptstyle T}}\psi(\tilde{z}_{j})\right)^{2}-\frac{2}{m}\sum_{i=1}^{m}\varphi(\tilde{x}_{i})^{\mathsf{\scriptscriptstyle T}}\psi(\tilde{z}_{i}).$ (8)
The spectral contrastive loss is the basis of the recent algorithm proposed in [34], where $\varphi$ and $\psi$ are parametrized by neural networks and optimized via stochastic gradient descent. While [34] motivate this objective heuristically, Eq. 6 shows that it directly targets the leading singular structure of $\mathcal{T}$. The use of objectives similar to Eq. 6 for operator learning or Eq. 7 for representation learning has a rich history that we detail in Appendix B.
Comparison to [34]. To justify the use of the contrastive loss in Eq. 8, [34] assume that the density $p(x,z)$ factorizes as
$p(x,z)=\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z),\quad\varphi:\mathcal{X}\to\mathbb{R}^{d},\quad\psi:\mathcal{Z}\to\mathbb{R}^{d}.$ (9)
We argue that this assumption is overly strong: it holds if and only if $\mathcal{T}$ has rank at most $d$, in which case the spectral features perfectly capture the structure of $\mathcal{T}$. This result is formalized in 5, Appendix A.3. We should instead interpret Eq. (9) as an approximation rather than a literal assumption. In practice, the learned features $\varphi,\psi$ provide a rank-$d$ approximation of $\mathcal{T}$, and the quality of this approximation governs the performance of the resulting Sieve 2SLS estimator.
Let $\hat{\varphi},\hat{\psi}$ be $d$-dimensional features trained with loss Eq. 8 with $\tilde{D}_{m}$ and define $\hat{\mathcal{T}}_{d}\doteq\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$. We make the following assumption:
Assumption 5.
$\{\hat{\psi}_{1},\ldots,\hat{\psi}_{d}\}$ and $\{\hat{\varphi}_{1},\ldots,\hat{\varphi}_{d}\}$ form linearly independent families.
This assumption is only made to simplify the discussion. If $\{\hat{\psi}_{1},\ldots,\hat{\psi}_{d}\}$ or $\{\hat{\varphi}_{1},\ldots,\hat{\varphi}_{d}\}$ is not linearly independent, one can replace $d$ in the following by the linear dimension of the family. Consider $\sigma_{d}>\sigma_{d+1}$ so that the unique minimizer of the population loss is given by $\mathcal{T}_{d}$. We quantify how the distance between $\hat{\mathcal{T}}_{d}$ and $\mathcal{T}_{d}$ affects the approximation error and the sieve measure of ill-posedness $\tau_{\hat{\varphi},d}$. We quantify how the distance between $\hat{\mathcal{T}}_{d}$ and $\mathcal{T}_{d}$ affects the generalization error of the 2SLS estimator.
Theorem 3.
Let Assumptions 1, 2, 4 and 5 hold and let $\hat{\varepsilon}_{d}\doteq\left\|\hat{\mathcal{T}}_{d}-\mathcal{T}_{d}\right\|_{\mathrm{op}}$ be such that $\hat{\varepsilon}_{d}<(1-1/\sqrt{2})\sigma_{d}$.
  • i)
    $\sigma_{d}^{-1}\leq\tau_{\hat{\varphi},d}\leq(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}$;
  • ii)
    Let $\hat{h}$ be the 2SLS estimator from Eq. 5, using features $\hat{\varphi}:\mathcal{X}\to\mathbb{R}^{d}$ and $\hat{\psi}:\mathcal{Z}\to\mathbb{R}^{d}$. Suppose $(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}\zeta_{\hat{\varphi},\hat{\psi},d}\sqrt{(\log d)/n}=o(1)$ and $\hat{\varepsilon}_{d}=o_{d}(\sigma_{d}^{2})$, then
    $\left\|\hat{h}-h_{0}\right\|_{L^{2}(X)}=O_{p}\left(\left\|(\Pi_{\mathcal{X},d})_{\perp}h_{0}\right\|_{L_{2}(X)}+\frac{\sqrt{2}\hat{\varepsilon}_{d}\|h_{0}\|_{L_{2}(X)}}{\sigma_{d}}+\frac{1}{\sigma_{d}-2\hat{\varepsilon}_{d}}\sqrt{\frac{d}{n}}\right)\,.$
We postpone the proof to Appendix A.3. The result quantifies how close the learned features must be to $\mathcal{T}_{d}$ in order to preserve the desirable properties of spectral 2SLS. Theorem 3-i) shows that as long as the learned operator $\hat{\mathcal{T}}_{d}$ is a good approximation of the true rank-$d$ truncation $\mathcal{T}_{d}$, the learned features inherit a favorable measure of ill-posedness. Theorem 3-ii) further ensures that if contrastive learning yields a near-optimal rank-$d$ approximation of $\mathcal{T}$, the downstream Sieve 2SLS estimator retains strong statistical guarantees. Importantly, the spectral features generalization error $\hat{\varepsilon}_{d}$ encapsulates the statistical and optimization errors in learning spectral features. It depends on the number of samples used to train the contrastive loss, the expressivity and architecture of the neural networks parametrizing $\varphi$ and $\psi$, and the complexity of the leading eigensubspaces of $\mathcal{T}$. As $\mathcal{L}_{d}(\varphi,\psi)$ converges to $\mathcal{L}_{d}^{\mathrm{min}}\doteq\left\|\mathcal{T}-\mathcal{T}_{d}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2},$ $\hat{\varepsilon}_{d}$ converges to 00 by Theorem 2. Precisely controlling $\hat{\varepsilon}_{d}$ remains an important question for future work.
Figure 1: Dependence of the MSE of instrumental variable regression on the decay rates of the spectrum and coefficients of the structural function. IV fitting and the MSE evaluations were repeated 500 times per parameter set, rendering the standard error of these estimates negligible.
5 Experiments
We evaluate the main theoretical insight of the paper: Sieve 2SLS with spectral features performs best when $h_{0}$ aligns with the top eigenspaces of $\mathcal{T}$ and its singular values decay slowly. Performance degrades with faster decay and weaker alignment. We design a synthetic NPIV setting where we control the conditional expectation operator $\mathcal{T}$ and its spectral decay and well as the alignment of $h_{0}$ with its eigenspaces. To simulate such a setting, we rely on the following procedure for generating samples from the NPIV model:
Proposition 3.
Let $\pi_{XZ}$ be a probability distribution on $\mathcal{X}\times\mathcal{Z}$ with marginals $\pi_{X}$ and $\pi_{Z}$. Let $\mathcal{T}:L_{2}(\mathcal{X},\pi_{X})\to L_{2}(\mathcal{Z},\pi_{Z})$ be the conditional expectation operator associated with $\pi_{XZ}$ and let $h_{0}\in L_{2}(\mathcal{X},\pi_{X})$. Sample $(X,Z)\sim\pi_{XZ}$, $V\sim\mathcal{N}(0,\sigma^{2})$ and set $U=\mathcal{T}h_{0}(Z)-h_{0}(X)+V$ and $Y=h_{0}(X)+U$. Then $(U,Z,X,Y)$ is a sample from the NPIV model of Eq. (1).
Let $d=11$, $\mathcal{X}=[0,2\pi]$, $\mathcal{Z}=[0,2\pi]$, $\sigma^{2}=0.1$, $f(x)=(\sin(x),\sin(2x),\ldots,\sin((d-1)x))$, $g(z)=(\sin(z),\sin(2z),\ldots,\sin((d-1)z))$, and let $P,Q$ be two orthogonal $(d-1)\times(d-1)$ matrices. Define $v:\mathcal{X}\to\mathbb{R}^{d-1}$ and $u:\mathcal{Z}\to\mathbb{R}^{d-1}$ as $v(x)=Pf(x)$ and $u(z)=Qg(z)$. We set $\mathcal{T}=\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=1}^{d-1}\sigma_{i}u_{i}\otimes v_{i}$. $(X,Z)$ is jointly sampled with rejection sampling. $Z,X$ are uniform on $[0,2\pi]$ and $X$ given $Z$ admits $\mathcal{T}$ as conditional expectation operator. We then sample $Y$ following 3. The multiplication of the trigonometric functions by the orthogonal matrices $P,Q$ ensures that the consecutive singular functions of $\mathcal{T}$ are comparably complex. By keeping the difficulty constant per singular function we can isolate the influence of the singular value decay on the performance of the final estimator. We set the singular values $\{\sigma_{i}\}_{i=1}^{d-1}$ to decay linearly, from some initial value $\sigma_{1}$ to a $\sigma_{d-1}=c_{\sigma}\sigma_{1}$, $c_{\sigma}\in[0,1]$. To control spectral alignment, we set $h_{0}=\sum_{i=1}^{d-1}\alpha_{i}v_{i}$ subject to $\|\alpha\|_{\ell_{2}}=1$, and allow $\{\alpha_{i}\}_{i=1}^{d-1}$ to decay linearly from $\alpha_{1}$ to $\alpha_{d-1}=c_{\alpha}\alpha_{1}$, $c_{\alpha}\in[0,1]$. Each value of $c_{\alpha}\in[0,1]$ controls the alignment of $h_{0}$ with the top singular functions of $\mathcal{T}$: larger values correspond to stronger alignment. Similarly, $c_{\sigma}\in[0,1]$ controls the ill-posedness of the problem: smaller values indicate faster decay of singular values and hence more severe ill-posedness.
Good/bad
Ugly
Very ugly
Figure 2: Comparison of the spectral methods’ performance depending on the alignment of $h_{0}$ with the singular functions of $\mathcal{T}$. Blue: True $h_{0}$; Orange: 2SLS estimate.
In Figure 1, we report the mean squared error (MSE) of two 2SLS estimators across a range of spectral decay rates ($c_{\sigma}$) and structural function decay rates ($c_{\alpha}$). Left: Oracle spectral features: The left panel uses 2SLS with fixed features $f$ and $g$ that form spectral features. This represents an oracle setting where feature learning is not needed. As expected, the MSE decreases from top to bottom as the problem becomes better conditioned (i.e., as $c_{\sigma}$ increases). Additionally, the MSE decreases from left to right, as smaller values of $c_{\alpha}$ correspond to stronger spectral alignment: more of $h_{0}$ is concentrated on the top singular functions of $\mathcal{T}$. These trends are fully consistent with the bound established in Corollary 1. Right: Learned features. In the right panel, we learn 50 features from data by minimizing the empirical contrastive loss (8) using a neural network. We then build a 2SLS estimator on top of the learned features. The same qualitative trends are observed: MSE decreases with slower singular value decay and better alignment. Moreover, when $c_{\sigma}$ is small (top rows), learning the entire eigenspaces becomes difficult. In this regime, performance is most sensitive to $c_{\alpha}$. In particular, if $h_{0}$ is concentrated on the top singular functions (small $c_{\alpha}$), good recovery is still possible. In contrast, when $c_{\sigma}$ is large (bottom rows), the singular functions are equally important and more easily learned, rendering the alignment of $h_{0}$ less critical. The mismatch between the oracle and learned settings is because we do not exactly recover the singular feature spaces; as predicted by Theorem 3. It is also worth noting that when $h_{0}$ is concentrated on high singular values, it may sometimes be beneficial for the singular values to decay quickly, by making it easier for the feature-learning model to learn the top singular functions well. This may explain the slight top-bottom increase in losses on the right end of the right panel.
The previous experiment assumes $h_{0}$ lies in the span of the singular functions of $\mathcal{T}$. Figure 2 illustrates three regimes. The left panel shows a well-aligned case (“good/bad”), where $h_{0}$ is entirely supported on the singular functions of $\mathcal{T}$, allowing for exact recovery as long as enough data is available. The middle panel represents a “partial alignment” regime, where $h_{0}$ has some overlap with the singular functions; in this case, recovery is imperfect, but the estimator can still capture some signal. Finally, the right panel illustrates the “misaligned” or “very ugly” regime, where $h_{0}$ is orthogonal to the singular functions. The instrument provides no information about the signal, and all spectral methods fail. We refer to Appendix C for further discussion of these scenarios and the experimental setup.
Appendix A Proofs
In this section we recall and prove the results of the main section.
A.1 Proofs of Section 2
We refer the reader to [4] for background on measure-theoretic notions used throughout this section. We assume that $\mathcal{X}$ and $\mathcal{Z}$ are standard Borel spaces, i.e., Borel subsets of complete separable metric spaces. This ensures the existence of regular conditional distributions and supports the disintegration of the joint law $\pi_{X,Z}$. Working directly with measures allows us to treat both continuous settings, where distributions admit densities with respect to the Lebesgue measure, and discrete or more general settings, where such densities may not exist.
To prove properties of the conditional expectation operator $\mathcal{T}$, it is helpful to express it in terms of a Markov kernel. We define a Markov kernel $p:\mathcal{Z}\times\mathcal{F}_{\mathcal{X}}\to[0,1]$ such that for any measurable set $A\in\mathcal{F}_{\mathcal{X}}$,
$p(z,A)=\mathbb{P}[X\in A\mid Z=z]\,,$
meaning $p(z,\cdot)$ is a regular conditional distribution of $X$ given $Z=z$. Then for all $h\in L_{2}(X)$,
[…]
By the disintegration theorem, for all measurable sets $A\subset\mathcal{X}$ and $B\subset\mathcal{Z}$, the joint distribution admits the disintegration
[…]
and therefore, the joint distribution $\pi_{X,Z}$ can be decomposed as
$d\pi_{X,Z}(x,z)=p(z,dx)\,d\pi_{Z}(z),$ (11)
where $p(z,dx)$ is a Markov kernel as defined above. This decomposition holds in general, even when $\pi_{X,Z}$ is not absolutely continuous with respect to $\pi_{X}\otimes\pi_{Z}$.
Under Assumption 2, the joint distribution $\pi_{X,Z}$ is absolutely continuous with respect to the product measure $\pi_{X}\otimes\pi_{Z}$, and thus admits a density
$p(x,z)\doteq\frac{d\pi_{X,Z}}{d(\pi_{X}\otimes\pi_{Z})}(x,z).$ (12)
In order to prove 1, we use the following lemma.
Lemma 1.
Under Assumption 2, for $\pi_{X}\otimes\pi_{Z}$-almost every $(x,z)$,
$p(x,z)=\frac{dp(z,\cdot)}{d\pi_{X}}(x).$
Proof.
From Eq. 11, we have
$d\pi_{X,Z}(x,z)=p(z,dx)\,d\pi_{Z}(z).$
On the other hand, by Eq. 12,
$d\pi_{X,Z}(x,z)=p(x,z)\,d\pi_{X}(x)\,d\pi_{Z}(z).$
Comparing both expressions and using the uniqueness of the Radon–Nikodym derivative, we conclude that for $\pi_{X}\otimes\pi_{Z}$-almost every $(x,z)$,
$p(z,dx)=p(x,z)\,d\pi_{X}(x),$
i.e., $p(z,\cdot)\ll\pi_{X}$, and
$\frac{dp(z,\cdot)}{d\pi_{X}}(x)=p(x,z).$
This concludes the proof. ∎
[…]
We recall and prove 1. This is a classical result, see, e.g., [25].
See 1
Proof.
Take $(e_{i})_{i\geq 1}$ any orthonormal basis (ONB) of $L_{2}(X)$. Then, using the definition of the Hilbert–Schmidt norm:
$\left\|\mathcal{T}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}$ $=\sum_{i\geq 1}\|\mathcal{T}e_{i}\|_{L_{2}(Z)}^{2}$ (definition of HS norm via ONB)
[…]
$=\int_{\mathcal{Z}}\sum_{i\geq 1}\langle e_{i},p(\cdot,z)\rangle_{L_{2}(X)}^{2}\,d\pi_{Z}(z)$ (Fubini’s theorem)
[…]
<<$\infty$< + ∞ (by Assumption 2, since $p\in L_{2}(\pi_{X}\otimes\pi_{Z})$).
Note that the proof shows in addition that $\left\|\mathcal{T}\right\|_{\scriptscriptstyle\mathrm{HS}}=\left\|p\right\|_{L_{2}\left(\pi_{X}\otimes\pi_{Z}\right)}$.
A.2 Proofs of Section 3
Proposition 4.
Let $\mathcal{H}_{\varphi,d}=\operatorname{span}\{\varphi_{1},\ldots,\varphi_{d}\}$, and let $\omega=\dim(\mathcal{H}_{\varphi,d})$. Then:
  1. i)
    $\tau_{\varphi,d}\geq\sigma_{\omega}^{-1}$.
  2. ii)
    If $\mathcal{H}_{\varphi,d}\subseteq\mathcal{V}_{\iota}=\operatorname{span}\{v_{1},\ldots,v_{\iota}\}$, then $\tau_{\varphi,d}\leq\sigma_{\iota}^{-1}$.
  3. iii)
    For any $d\geq 1$, the minimal value $\tau_{\varphi,d}=\sigma_{d}^{-1}$ is achieved when $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$.
4 is due to Lemma 1 of [3] and 2 corresponds to part (iii). We provide a full proof for completeness.
Proof.
We first prove (i). Since $\dim(\mathcal{H}_{\varphi,d})=\omega$, and the subspace $\mathcal{V}_{\omega-1}$ has dimension $\omega-1$, there exists $\tilde{h}\in\mathcal{H}_{\varphi,d}\cap\mathcal{V}_{\omega-1}^{\perp}$ with $\|\tilde{h}\|_{L_{2}(X)}=1$. Then,
$\tau_{\varphi,d}^{-2}$ $=\inf_{\begin{subarray}{c}h\in\mathcal{H}_{\varphi,d}\\ \|h\|_{L_{2}(X)}=1\end{subarray}}\|\mathcal{T}h\|_{L_{2}(Z)}^{2}$
$\leq\|\mathcal{T}\tilde{h}\|_{L_{2}(Z)}^{2}$
$\leq\sup_{\begin{subarray}{c}h\in\mathcal{V}_{\omega-1}^{\perp}\\ \|h\|_{L_{2}(X)}=1\end{subarray}}\|\mathcal{T}h\|_{L_{2}(Z)}^{2}$
$=\sup_{\begin{subarray}{c}h\in\mathcal{V}_{\omega-1}^{\perp}\\ \|h\|=1\end{subarray}}\langle\mathcal{T}^{*}\mathcal{T}h,h\rangle_{L_{2}(X)}=\sigma_{\omega}^{2},$
where the last line follows from the min–max theorem for compact self-adjoint operators.
We now prove (ii). If $h\in\mathcal{H}_{\varphi,d}\subseteq\mathcal{V}_{\iota}$ and $\|h\|_{L_{2}(X)}=1$, then
$h=\sum_{i=1}^{\iota}\langle h,v_{i}\rangle v_{i},$
so by orthonormality of $(u_{i})$,
$\|\mathcal{T}h\|_{L_{2}(Z)}^{2}=\left\|\sum_{i=1}^{\iota}\sigma_{i}\langle h,v_{i}\rangle u_{i}\right\|_{L_{2}(Z)}^{2}=\sum_{i=1}^{\iota}\sigma_{i}^{2}\langle h,v_{i}\rangle^{2}\geq\sigma_{\iota}^{2}\sum_{i=1}^{\iota}\langle h,v_{i}\rangle^{2}=\sigma_{\iota}^{2}.$
Taking the infimum over unit-norm $h\in\mathcal{H}_{\varphi,d}$ gives $\tau_{\varphi,d}^{-2}\geq\sigma_{\iota}^{2}$, i.e., $\tau_{\varphi,d}\leq\sigma_{\iota}^{-1}$.
Finally, (iii) follows by taking $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$, so that $\omega=d=\iota$, and both bounds in (i) and (ii) match. ∎
We recall and prove Corollary 1:
See 1
Proof.
Let us verify that Assumption 3 holds under Definition 2, i.e., when $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$ and $\mathcal{H}_{\psi,d}=\mathcal{U}_{d}$.
(i) Since the image of $\mathcal{V}_{d}$ under $\mathcal{T}$ lies in $\mathcal{U}_{d}$, we have
$(\Pi_{\psi,d})_{\perp}\mathcal{T}h=0\quad\text{for all }h\in\mathcal{H}_{\varphi,d}.$
Hence,
$\sup_{\begin{subarray}{c}h\in\mathcal{H}_{\varphi,d}\\ \|h\|_{L_{2}}=1\end{subarray}}\left\|(\Pi_{\psi,d})_{\perp}\mathcal{T}h\right\|_{L^{2}(Z)}=0=o\left(\tau_{\varphi,d}^{-1}\right),$
and condition (i) is satisfied.
(ii) Let $h_{0}\in L_{2}(X)$, and consider its projection onto the orthogonal complement of $\mathcal{H}_{\varphi,d}$. Since $\mathcal{T}$ is compact with singular value decomposition $\mathcal{T}=\sum_{i=1}^{\infty}\sigma_{i}u_{i}\otimes v_{i}$ under Assumption 2, we have
$\|\mathcal{T}(\Pi_{\varphi,d})_{\perp}h_{0}\|_{L_{2}(Z)}^{2}=\sum_{i>d}\sigma_{i}^{2}\langle h_{0},v_{i}\rangle^{2}\leq\sigma_{d}^{2}\sum_{i>d}\langle h_{0},v_{i}\rangle^{2}=\sigma_{d}^{2}\|(\Pi_{\varphi,d})_{\perp}h_{0}\|_{L_{2}(X)}^{2}.$
This implies, using that $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by 2,
$\|\mathcal{T}(\Pi_{\varphi,d})_{\perp}h_{0}\|_{L_{2}(Z)}\leq\sigma_{d}\|(\Pi_{\varphi,d})_{\perp}h_{0}\|_{L_{2}(X)}=\tau_{\varphi,d}^{-1}\|(\Pi_{\varphi,d})_{\perp}h_{0}\|_{L_{2}(X)},$
so Assumption 3-(ii) holds with constant $C=1$.
Since both parts of Assumption 3 hold, and $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by 2, we may apply Theorem 1, which yields the desired result. ∎
A.3 Proofs of Section 4
The following result shows that Eq. 9 holds if and only if $\operatorname{rank}(\mathcal{T})\leq d$.
Proposition 5.
Let $\varphi:\mathcal{X}\to\mathbb{R}^{d}$, $\psi:\mathcal{Z}\to\mathbb{R}^{d}$ with components $\varphi_{i}\in L_{2}(X)$, $\psi_{i}\in L_{2}(Z)$. Then
$p(x,z)=\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)\quad\text{if and onlyif}\quad\mathcal{T}=\sum_{i=1}^{d}\psi_{i}\otimes\varphi_{i}.$
Proof.
$\Rightarrow$ Assume $p(x,z)=\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)$. Then, by Eq. 10, for all $h\in L_{2}(X)$,
[…]
$=\sum_{i=1}^{d}\psi_{i}(z)\langle h,\varphi_{i}\rangle_{L_{2}(X)}=\left(\sum_{i=1}^{d}\psi_{i}\otimes\varphi_{i}\right)(h)(z),$
so $\mathcal{T}=\sum_{i=1}^{d}\psi_{i}\otimes\varphi_{i}$.
$\Leftarrow$ Now assume $\mathcal{T}=\sum_{i=1}^{d}\psi_{i}\otimes\varphi_{i}$. Let $f\in L_{2}(\pi_{X}\otimes\pi_{Z})$. We want to show:
$\int f(x,z)p(x,z)\,d(\pi_{X}\otimes\pi_{Z})(x,z)=\int f(x,z)\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)\,d(\pi_{X}\otimes\pi_{Z})(x,z).$
Starting with the right-hand side:
[…]
Since both integrals agree for all $f\in L_{2}(\pi_{X}\otimes\pi_{Z})$, and $p(x,z)$ is the Radon–Nikodym derivative of $\pi_{X,Z}$ with respect to $\pi_{X}\otimes\pi_{Z}$, it follows by uniqueness that
$p(x,z)=\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)\quad\text{for }(\pi_{X}\otimes\pi_{Z})\text{-almost every }(x,z).$
In order to prove Theorem 3 we prove intermediate results.
Proposition 6.
Assumption 5 holds if and only if $\operatorname{rank}(\hat{\mathcal{T}}_{d})=d$.
Proof.
Assume that Assumption 5 holds, i.e., $\{\hat{\varphi}_{1},\ldots,\hat{\varphi}_{d}\}$ and $\{\hat{\psi}_{1},\ldots,\hat{\psi}_{d}\}$ are linearly independent. We show that $\operatorname{rank}(\hat{\mathcal{T}}_{d})=d$.
Since $\hat{\mathcal{T}}_{d}h=\sum_{i=1}^{d}\langle h,\hat{\varphi}_{i}\rangle\hat{\psi}_{i}$, we can write $\hat{\mathcal{T}}_{d}=A\circ\Phi$, where:
  • $\Phi:L_{2}(X)\to\mathbb{R}^{d}$, defined by $h\mapsto(\langle h,\hat{\varphi}_{1}\rangle,\ldots,\langle h,\hat{\varphi}_{d}\rangle)$,
  • $A:\mathbb{R}^{d}\to L_{2}(Z)$, defined by $a\mapsto\sum_{i=1}^{d}a_{i}\hat{\psi}_{i}$.
Under Assumption 5, the family $\{\hat{\varphi}_{i}\}$ is linearly independent, so $\Phi$ is surjective. Similarly, the family $\{\hat{\psi}_{i}\}$ is linearly independent, so $A$ is injective. Therefore, the composition $A\circ\Phi$ has image of dimension
$\operatorname{rank}(\hat{\mathcal{T}}_{d})=\dim(\mathcal{R}(A\circ\Phi))=\dim(\mathcal{R}(A))=d,$
where the second equality follows from the surjectivity of $\Phi$, and the third from the injectivity of $A$.
Conversely, suppose $\operatorname{rank}(\hat{\mathcal{T}}_{d})=d$. Then $\hat{\mathcal{T}}_{d}$ has image of dimension $d$, which implies that the $\hat{\psi}_{i}$ must be linearly independent. The same reasoning applied to $\mathcal{T}^{*}$ shows that $\{\hat{\varphi}_{1},\ldots,\hat{\varphi}_{d}\}$ is linearly independent.
Under Assumption 5, we can therefore write $\hat{\mathcal{T}}_{d}$ in the following SVD form:
$\hat{\mathcal{T}}_{d}=\sum_{i=1}^{d}\hat{\sigma}_{i}\hat{u}_{i}\otimes\hat{v}_{i},$ (14)
where $\hat{\sigma}_{1}\geq\hat{\sigma}_{2}\geq\cdots\geq\hat{\sigma}_{d}>0$ are the singular values of $\hat{\mathcal{T}}$ and $\hat{u}_{i}$ and $\hat{v}_{i}$ are the left and right singular functions of $\hat{\mathcal{T}}$.
Proposition 7.
Under Assumption 5, $\mathcal{H}_{\hat{\varphi},d}=\operatorname{span}\left\{\hat{v}_{1},\ldots,\hat{v}_{d}\right\}$ and $\mathcal{H}_{\hat{\psi},d}=\operatorname{span}\left\{\hat{u}_{1},\ldots,\hat{u}_{d}\right\}$.
Proof.
We have by definition of the SVD that
$\mathcal{N}(\hat{\mathcal{T}}_{d})^{\perp}=\operatorname{span}\left\{\hat{v}_{1},\ldots,\hat{v}_{d}\right\}.$
To show that $\mathcal{H}_{\hat{\varphi},d}=\operatorname{span}\left\{\hat{v}_{1},\ldots,\hat{v}_{d}\right\}$, it suffices to prove that
$\mathcal{N}(\hat{\mathcal{T}}_{d})=\operatorname{span}\left\{\hat{\varphi}_{1},\ldots,\hat{\varphi}_{d}\right\}^{\perp}.$
Let $h\in L_{2}(X)$. Then:
$h\in\mathcal{N}(\hat{\mathcal{T}}_{d})$ $\iff\hat{\mathcal{T}}_{d}h=0=\sum_{i=1}^{d}\langle h,\hat{\varphi}_{i}\rangle\hat{\psi}_{i}$
$\iff\sum_{i=1}^{d}\langle h,\hat{\varphi}_{i}\rangle\hat{\psi}_{i}=0$
$\iff\langle h,\hat{\varphi}_{i}\rangle=0\quad\forall i\in[d],$
where the last equivalence follows from the linear independence of $\{\hat{\psi}_{i}\}_{i=1}^{d}$.
Therefore, $\mathcal{N}(\hat{\mathcal{T}}_{d})=\operatorname{span}\{\hat{\varphi}_{1},\ldots,\hat{\varphi}_{d}\}^{\perp}$, which implies
$\mathcal{H}_{\hat{\varphi},d}=\mathcal{N}(\hat{\mathcal{T}}_{d})^{\perp}=\operatorname{span}\left\{\hat{v}_{1},\ldots,\hat{v}_{d}\right\}.$
The proof for $\mathcal{H}_{\hat{\psi},d}=\operatorname{span}\left\{\hat{u}_{1},\ldots,\hat{u}_{d}\right\}$ is analogous, using the adjoint $\hat{\mathcal{T}}_{d}^{*}$. ∎
We now recall and prove Theorem 3: See 3
Proof.
We first show that $\sigma_{d}^{-1}\leq\tau_{\hat{\varphi},d}\leq(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}$. First note that if $h\in\mathcal{H}_{\hat{\varphi},d}$, then by Proposition 7, $h\in\mathcal{V}_{d}$. Using the SVD of $\hat{\mathcal{T}}_{d}$ from Eq. 14, and writing $h=\sum_{i=1}^{d}\langle h,\hat{v}_{i}\rangle\hat{v}_{i}$, we obtain:
$\hat{\mathcal{T}}_{d}h=\sum_{i=1}^{d}\hat{\sigma}_{i}\langle h,\hat{v}_{i}\rangle\hat{u}_{i},\quad\text{so}\quad\|\hat{\mathcal{T}}_{d}h\|_{L_{2}(Z)}^{2}=\sum_{i=1}^{d}\hat{\sigma}_{i}^{2}\langle h,\hat{v}_{i}\rangle^{2}\geq\hat{\sigma}_{d}^{2}\|h\|_{L_{2}(X)}^{2}.$
Next, observe that for all $h\in L_{2}(X)$,
$\|\mathcal{T}h\|_{L_{2}(Z)}\geq\|\mathcal{T}_{d}h\|_{L_{2}(Z)}.$
By the reverse triangle inequality, the bound $\left\|\hat{\mathcal{T}}_{d}-\mathcal{T}_{d}\right\|_{\mathrm{op}}\leq\hat{\varepsilon}_{d}$, and Weyl’s inequality for singular values, we have, for all $h\in L_{2}(X)$, with $\|h\|_{L_{2}(X)}=1$:
$\|\mathcal{T}h\|_{L_{2}(Z)}$ $\geq\|\mathcal{T}_{d}h\|_{L_{2}(Z)}$
$\geq\|\hat{\mathcal{T}}_{d}h\|_{L_{2}(Z)}-\left\|\hat{\mathcal{T}}_{d}-\mathcal{T}_{d}\right\|_{\mathrm{op}}\cdot\|h\|_{L_{2}(X)}$
$\geq(\hat{\sigma}_{d}-\hat{\varepsilon}_{d})\|h\|_{L_{2}(X)}$
$=\hat{\sigma}_{d}-\hat{\varepsilon}_{d}$
$\geq\sigma_{d}-2\hat{\varepsilon}_{d}.$
This shows that
$\tau_{\hat{\varphi},d}^{-1}\geq\sigma_{d}-2\hat{\varepsilon}_{d}\quad\Rightarrow\quad\tau_{\hat{\varphi},d}\leq(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1},$
which holds under the assumption $\hat{\varepsilon}_{d}<(1-1/\sqrt{2})\sigma_{d}<\sigma_{d}/2$. On the other hand, since $\dim(\mathcal{H}_{\hat{\varphi},d})=d$ by Assumption 5, Proposition 4-(i) implies
$\sigma_{d}\quad\Rightarrow\quad\tau_{\hat{\varphi},d}\geq\sigma_{d}^{-1}.$
Next, we show how to bound the sieve approximation error, write:
$\left\|(\Pi_{\hat{\varphi},d})_{\perp}h_{0}\right\|_{L_{2}(X)}$ $=\left\|(\Pi_{\hat{\varphi},d})_{\perp}\left((\Pi_{\mathcal{X},d})_{\perp}h_{0}+\Pi_{\mathcal{X},d}h_{0}\right)\right\|_{L_{2}(X)}$
$\leq\left\|(\Pi_{\mathcal{X},d})_{\perp}h_{0}\right\|_{L_{2}(X)}+\left\|(\Pi_{\hat{\varphi},d})_{\perp}\Pi_{\mathcal{X},d}h_{0}\right\|_{L_{2}(X)}$
$\leq\left\|(\Pi_{\mathcal{X},d})_{\perp}h_{0}\right\|_{L_{2}(X)}+\left\|(\Pi_{\hat{\varphi},d})_{\perp}\Pi_{\mathcal{X},d}\right\|_{\mathrm{op}}\cdot\|h_{0}\|_{L_{2}(X)}.$
By Wedin’s sin–$\Theta$ theorem (Theorem 2.9 and Eq. (2.26a) in [8]), and under the assumption $\hat{\varepsilon}_{d}<(1-1/\sqrt{2})\sigma_{d}$, we have:
$\left\|(\Pi_{\hat{\varphi},d})_{\perp}\Pi_{\mathcal{X},d}\right\|_{\mathrm{op}}\leq\frac{\sqrt{2}\hat{\varepsilon}_{d}}{\sigma_{d}}.$
Therefore,
$\left\|(\Pi_{\hat{\varphi},d})_{\perp}h_{0}\right\|_{L_{2}(X)}\leq\left\|(\Pi_{\mathcal{X},d})_{\perp}h_{0}\right\|_{L_{2}(X)}+\frac{\sqrt{2}\hat{\varepsilon}_{d}\|h_{0}\|_{L_{2}(X)}}{\sigma_{d}},$
To conclude we now check Assumption 3. Let us introduce
$s_{\hat{\psi},\hat{\varphi},d}^{-1}=\sup_{h\in\mathcal{H}_{\hat{\varphi},d},~{}h\neq 0}\frac{\|h\|_{L_{2}(X)}}{\|\Pi_{\hat{\psi},d}\mathcal{T}h\|_{L_{2}(Z)}}\geq\tau_{\hat{\varphi},d}.$
The unique use of Assumption 3-i) in the proof of Theorem B.1 [6] is to prove that $s_{\hat{\psi},\hat{\varphi},d}^{-1}=O(\tau_{\hat{\varphi},d})$ as $d\to+\infty$, which we now directly prove. First, note that for any $h\in L_{2}(X),$
$\left\|\mathcal{T}\Pi_{\hat{\varphi},d}h\right\|_{L_{2}(Z)}$ $=\left\|\mathcal{T}\left(\Pi_{\hat{\varphi},d}+\Pi_{\mathcal{X},d}-\Pi_{\mathcal{X},d}\right)h\right\|_{L_{2}(Z)}$
$\leq\left\|\mathcal{T}_{d}h\right\|_{L_{2}(Z)}+\left\|\Pi_{\hat{\varphi},d}-\Pi_{\mathcal{X},d}\right\|_{\mathrm{op}}\left\|h\right\|_{L_{2}(X)}$
$\leq\left\|\mathcal{T}_{d}h\right\|_{L_{2}(Z)}+\frac{\sqrt{2}\hat{\varepsilon}_{d}\|h\|_{L_{2}(X)}}{\sigma_{d}},$
where we used $\left\|\mathcal{T}\right\|_{\mathrm{op}}\leq 1$, the definition of $\mathcal{T}_{d}$ and Wedin’s sin–$\Theta$ theorem. On the other hand,
$\left\|\Pi_{\hat{\psi},d}\mathcal{T}\Pi_{\hat{\varphi},d}h\right\|_{L_{2}(Z)}$ $\geq\left\|\Pi_{\hat{\psi},d}\mathcal{T}\Pi_{\varphi,d}h\right\|_{L_{2}(Z)}-\left\|\Pi_{\hat{\psi},d}\mathcal{T}\left(\Pi_{\hat{\varphi},d}-\Pi_{\mathcal{X},d}\right)h\right\|_{L_{2}(Z)}$
$\geq\left\|\Pi_{\psi,d}\mathcal{T}\Pi_{\varphi,d}h\right\|_{L_{2}(Z)}-\left\|(\Pi_{\psi,d}-\Pi_{\hat{\psi},d})\mathcal{T}\Pi_{\varphi,d}h\right\|_{L_{2}(Z)}-\frac{\sqrt{2}\hat{\varepsilon}_{d}\|h\|_{L_{2}(X)}}{\sigma_{d}}$
$\geq\left\|\mathcal{T}_{d}h\right\|_{L_{2}(Z)}-\frac{2\sqrt{2}\hat{\varepsilon}_{d}\|h\|_{L_{2}(X)}}{\sigma_{d}}.$
We therefore obtain that,
$s_{\hat{\psi},\hat{\varphi},d}^{-1}\leq\sup_{h\in\mathcal{H}_{\hat{\varphi},d},~{}h\neq 0}\frac{1}{\frac{\left\|\mathcal{T}\Pi_{\hat{\varphi},d}h\right\|_{L_{2}(Z)}}{\|h\|_{L_{2}(X)}}-\frac{3\sqrt{2}\hat{\varepsilon}_{d}}{\sigma_{d}}}\leq\tau_{\hat{\varphi},d}\times\frac{1}{1-\frac{3\sqrt{2}\hat{\varepsilon}_{d}\tau_{\hat{\varphi},d}}{\sigma_{d}}},$
where the inequality is valid for $d$ large enough as long as $3\sqrt{2}\hat{\varepsilon}_{d}\tau_{\hat{\varphi},d}\sigma_{d}^{-1}=o(1)$. Re-using that $\tau_{\hat{\varphi},d}\leq(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}$, a sufficient condition is
$3\sqrt{2}\frac{\hat{\varepsilon}_{d}}{\sigma_{d}}\frac{1}{\sigma_{d}-2\hat{\varepsilon}_{d}}=o(1),$
which is satisfied if $\hat{\varepsilon}_{d}=o(\sigma_{d}^{2})$. We therefore conclude that if $\hat{\varepsilon}_{d}=o(\sigma_{d}^{2})$ then $s_{\hat{\psi},\hat{\varphi},d}^{-1}=O(\tau_{\hat{\varphi},d})$. Finally, we check Assumption 3-ii). First note that by definition of $\Pi_{\hat{\varphi},d}$ and $\hat{\mathcal{T}}_{d}$, we have $\hat{\mathcal{T}}_{d}(\Pi_{\hat{\varphi},d})_{\perp}=0$, therefore,
$\tau_{\hat{\varphi},d}\left\|\mathcal{T}(\Pi_{\hat{\varphi},d})_{\perp}h_{0}\right\|$ $=\tau_{\hat{\varphi},d}\left\|(\mathcal{T}-\hat{\mathcal{T}}_{d})(\Pi_{\hat{\varphi},d})_{\perp}h_{0}\right\|$
$\leq\tau_{\hat{\varphi},d}\left(\left\|(\mathcal{T}-\mathcal{T}_{d})(\Pi_{\hat{\varphi},d})_{\perp}h_{0}\right\|+\left\|(\mathcal{T}_{d}-\hat{\mathcal{T}}_{d})(\Pi_{\hat{\varphi},d})_{\perp}h_{0}\right\|\right)$
$\leq\tau_{\hat{\varphi},d}(\sigma_{d+1}+\hat{\varepsilon}_{d})\left\|(\Pi_{\hat{\varphi},d})_{\perp}h_{0}\right\|$
$\leq\frac{\sigma_{d+1}+\hat{\varepsilon}_{d}}{\sigma_{d}-2\hat{\varepsilon}_{d}}\left\|(\Pi_{\hat{\varphi},d})_{\perp}h_{0}\right\|,$
We conclude using that $\hat{\varepsilon}_{d}=o(\sigma_{d}^{2})$ implies $\hat{\varepsilon}_{d}=o(\sigma_{d})$ which implies $(\sigma_{d+1}+\hat{\varepsilon}_{d})(\sigma_{d}-2\hat{\varepsilon}_{d})=O(1)$. ∎
We now prove the equivalence between Eq. 6 and Eq. 7. While the proof strategy follows that of Kostic et al. [24, Theorem 1], we include our own version for completeness, as the parametrization of the learned operator differs. Specifically, [24] directly approximate the truncated operator $\mathcal{T}_{d}$ using a singular value decomposition of the form $\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=2}^{d}\hat{\sigma}_{i}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$, where the singular values $\hat{\sigma}_{i}$ are learned explicitly via a separate network. In contrast, our approach learns only the feature maps $(\varphi_{i},\psi_{i})$ and defines the approximation $\mathcal{T}_{d}(\hat{\varphi},\hat{\psi})=\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$. As such, we provide a short self-contained derivation adapted to our setting for clarity. Recall that the Eckart–Young–Mirsky formulation of the objective is defined as
$\mathcal{L}_{d}(\varphi,\psi)=\left\|\mathcal{T}_{d}(\varphi,\psi)-\mathcal{T}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}=\left\|\mathcal{T}_{d}(\varphi,\psi)\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}-2\langle\mathcal{T}_{d}(\varphi,\psi),\mathcal{T}\rangle_{\scriptscriptstyle\mathrm{HS}}+\left\|\mathcal{T}_{d}(\varphi,\psi)\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}\,.$
Proposition 8.
It holds that,
$\left\|\mathcal{T}_{d}(\varphi,\psi)\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}-2\langle\mathcal{T}_{d}(\varphi,\psi),\mathcal{T}\rangle_{\scriptscriptstyle\mathrm{HS}}=\mathbb{E}_{X}\mathbb{E}_{Z}\left[\left(\varphi(X)^{\mathsf{\scriptscriptstyle T}}\psi(Z)\right)^{2}\right]-2\mathbb{E}_{X,Z}\left[\varphi(X)^{\mathsf{\scriptscriptstyle T}}\psi(Z)\right].$
Proof.
Let us introduce
$\Phi:\mathbb{R}^{d}\to L_{2}(X),\quad c\mapsto\sum_{i=1}^{d}c_{i}\varphi_{i},\qquad\Psi:\mathbb{R}^{d}\to L_{2}(Z),\quad c\mapsto\sum_{i=1}^{d}c_{i}\psi_{i},$
such that $\mathcal{T}_{d}(\varphi,\psi)=\Psi\Phi^{*}$, $\Psi^{*}\Psi=\mathbb{E}[\psi(Z)\psi(Z)^{\mathsf{\scriptscriptstyle T}}]$ and $\Phi^{*}\Phi=\mathbb{E}[\varphi(X)\varphi(X)^{\mathsf{\scriptscriptstyle T}}]$. On one hand, we have:
$\left\|\mathcal{T}_{d}(\varphi,\psi)\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}$ $=\left\|\Psi\Phi^{*}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}$
$=\operatorname{Tr}\left(\Psi^{*}\Psi\Phi^{*}\Phi\right)$
$=\operatorname{Tr}\left(\mathbb{E}[\psi(Z)\psi(Z)^{\mathsf{\scriptscriptstyle T}}]\mathbb{E}[\varphi(X)\varphi(X)^{\mathsf{\scriptscriptstyle T}}]\right)$
$=\mathbb{E}_{X}\mathbb{E}_{Z}\left[\operatorname{Tr}\left(\psi(Z)\psi(Z)^{\mathsf{\scriptscriptstyle T}}\varphi(X)\varphi(X)^{\mathsf{\scriptscriptstyle T}}\right)\right]$
$=\mathbb{E}_{X}\mathbb{E}_{Z}\left[\left(\psi(Z)^{\mathsf{\scriptscriptstyle T}}\varphi(X)\right)^{2}\right].$
On the other hand, we have:
$\langle\mathcal{T}_{d}(\varphi,\psi),\mathcal{T}\rangle_{\scriptscriptstyle\mathrm{HS}}$ $=\operatorname{Tr}\left(\Psi^{*}\mathcal{T}\Phi\right)$
$=\operatorname{Tr}\left(\mathbb{E}\left[\psi(Z)\mathbb{E}\left[\varphi(X)\mid Z\right]^{\mathsf{\scriptscriptstyle T}}\right]\right)$
$=\mathbb{E}_{XZ}\left[\operatorname{Tr}\left(\psi(Z)\varphi(X)^{\mathsf{\scriptscriptstyle T}}\right)\right]$
$=\mathbb{E}_{XZ}\left[\psi(Z)^{\mathsf{\scriptscriptstyle T}}\varphi(X)\right],$
which conclude the proof. ∎
As noted in [24], the above result does not require $\mathcal{T}$ to be Hilbert-Schmidt, or even compact. Indeed as $\mathcal{T}_{d}(\varphi,\psi)$ is a finite rank operator, $\langle\mathcal{T}_{d}(\varphi,\psi),\mathcal{T}\rangle_{\scriptscriptstyle\mathrm{HS}}$ is always well defined.
Appendix B Extended Related Work
As mentioned in Section 1, a central challenge in NPIV estimation is that it constitutes an ill-posed inverse problem. The literature has evolved along several methodological lines to address it. Two broad classes of estimation strategies have become particularly prominent:
  • Two-Stage Least-Squares (2SLS). This classical approach and its various generalizations involves a sequential estimation procedure. Typically, the first stage estimates the conditional expectation of the endogenous variable (or its features) given the instrument, and the second stage uses these predictions to estimate the structural function.
  • Saddle-point optimization. These methods, often rooted in a generalized method of moments (GMM) framework or duality principles, reformulate the NPIV estimation as finding an equilibrium in a min-max game. This often involves optimizing an objective function over a hypothesis space for the structural function and a test function space for moment conditions.
These two methodologies developed with distinct focuses. The 2SLS methods, with roots in classical econometrics [30, 10], have progressively incorporated more sophisticated nonparametric techniques. Saddle-point methods, on the other hand, have gained traction with the rise of machine learning, offering powerful tools for handling complex optimization problems arising from conditional moment restrictions [12, 28, 2].
B.0.0.0.1 2SLS approaches.
Early and influential nonparametric extensions of 2SLS employed sieve or series estimators. These methods approximate the unknown functions using a finite linear combination of basis functions, such as polynomials, splines, or Fourier series. The number of basis functions, or the “sieve dimension”, is allowed to increase with the sample size, enabling consistent estimation of the nonparametric functions.
A seminal contribution, by [30], provided identification results and a consistent nonparametric estimator for conditional moment restrictions. Their proposed NPIV estimator is an analog of 2SLS, where the first stage involves nonparametric estimation of conditional means of basis functions of $X$ given $Z$, and the second stage uses a series approximation for the structural function based on these first-stage predictions. Regularization is achieved by controlling the number of terms in the series approximation.
Building on similar principles, [17] proposed nonparametric methods based on both kernel444”Kernel” is meant here in the sense of density estimation, and not reproducing kernel: see [17, eq. (2.4)]. techniques and orthogonal series for estimating regression functions with instrumental variables. For their orthogonal series estimator, they transformed the marginal distributions of $X$ and $Z$ to be uniform and used Fourier expansions. The estimated Fourier coefficients of the structural function were obtained via a regularized regression involving estimated coefficients from the first stage. They were among the first to derive optimal convergence rates for this class of problems, explicitly linking these rates to the “difficulty” of the ill-posed inverse problem, which is characterized by the eigenvalues of the underlying integral operator.
[10] proposed an estimation procedure based on Tikhonov regularization. This involves regularizing the inverse of the integral operator $\mathcal{T}$ (or its empirical counterpart) to stabilize the solution. Specifically, denoting $\hat{\mathcal{T}}$ and $\hat{r}_{0}$ empirical estimates (in their case, computed with kernel density estimators), the Tikhonov regularized solution is of the form $h_{\alpha}=\left(\hat{\mathcal{T}}^{\star}\hat{\mathcal{T}}+\alpha I\right)^{-1}\hat{\mathcal{T}}^{\star}\hat{r}_{0}$, where $\alpha>0$. They presented asymptotic properties of their estimator, including consistency and convergence rates, which depend on the smoothness of the structural function and the degree of ill-posedness of $\mathcal{T}$.
[33] introduced Kernel Instrumental Variable Regression (KIV), a direct nonparametric generalization of 2SLS. KIV models the relationships between the different variables as nonlinear functions in reproducing kernel Hilbert spaces (RKHSs). In stage 1, it learns a conditional mean embedding $\mu(z)=\mathbb{E}\left[\varphi(X)\mathrel{}\middle|\mathrel{}Z=z\right]$ where $\varphi(X)$ represents features of $X$ mapped into an RKHS $\mathcal{H}_{X}$. This learning is framed as a vector-valued kernel ridge regression, effectively estimating the conditional expectation operator $E:\mathcal{H}_{X}\to\mathcal{H}_{Z}$. In stage 2, a scalar-valued kernel ridge regression of the outcome $Y$ on the estimated means $\hat{\mu}(Z)$ is performed to estimate the structural function. The authors prove the consistency of KIV in the projected norm under mild conditions, and derive conditions under which KIV achieves minimax optimal convergence rates. The analysis was later improved by [29], who establish convergence in $L_{2}$ norm rather than the projected norm.
B.0.0.0.2 Saddle-point approaches.
These arise from reformulating the conditional moment restrictions that define the NPIV problem (Eq. 1), frequently conceptualized as a zero-sum game. One player selects a function $h$ from a hypothesis space $\mathcal{H}$ to minimize the objective. The other player, the “adversary” or “witness”, selects a test function $g$ from a test function space $\mathcal{G}$ to maximize the objective. The adversary’s role is to select the function that maximizes the violation of the moment condition by the current choice of $h$. A notable advantage of this formulation is that it can allow bypassing the direct estimation of conditional expectations. Saddle-point objectives in NPIV regression can be derived in different ways.
One can frame the NPIV problem as finding a solution $h_{0}$ that minimizes a certain criterion (e.g., a norm) subject to satisfying the moment condition $\mathcal{T}h_{0}=r_{0}$. The Lagrangian of this constrained problem then leads to a minimax objective. [2] target the least-norm solution
$h_{0}=\operatorname*{arg\,min}_{h\in L_{2}(X)}\frac{1}{2}\left\|h\right\|_{L_{2}(X)}^{2}\quad\text{subject to}\quad\mathcal{T}h=r_{0}\,.$
The corresponding Lagrangian is given by $L\left(f,g\right)=\frac{1}{2}\left\|h\right\|_{L_{2}(X)}^{2}+\left\langle r_{0}-\mathcal{T}h,g\right\rangle_{L_{2}(Z)}$ for $g\in L_{2}(Z)$. Their method achieves strong $L_{2}$ error rates under a source condition and realizability assumptions. Notably, their approach does not require the often-problematic closedness condition or uniqueness of the IV solution. [28] consider a similar approach, focusing on the setting where the function classes are formed by neural networks, but use different techniques to analyze their method (online learning and neural tangent kernel theory).
Another approach to deriving a saddle-point problem involves using Fenchel duality to transform a squared error loss involving conditional expectations. This is the approach taken by [34], which we study in this paper.
A third method is to directly use the unconditional moment formulation $\mathbb{E}\left[\left(Y-h(X)\right)g(Z)\right]=0$. [12] define their criterion as the maximum moment deviation,
$h_{0}=\operatorname*{arg\,inf}_{h\in\mathcal{H}}\sup_{g\in\mathcal{G}}\mathbb{E}\left[\left(Y-h(X)\right)g(Z)\right]=\Psi\left(h,g\right)\,,$
and define the estimator $\hat{h}=\operatorname*{arg\,min}_{h}\sup_{g}\mathbb{E}_{n}\left[\left(Y-h(X)\right)g(Z)\right]+\mu R_{1}(h)-\lambda R_{2}(g)$, where $R_{1}$, $R_{2}$ are regularizers. Their key theoretical result is that the statistical estimation rate, in terms of projected mean squared error $\|\mathcal{T}(\hat{h}-h_{0})\|_{2}$, scales with the critical radii of the hypothesis space $\mathcal{H}$ and the test function space $\mathcal{G}$. This holds under some closedness assumption, namely that $\mathbb{E}\left[h(X)-h^{\prime}(X)\mathrel{}\middle|\mathrel{}Z\right]\in\mathcal{G}$ for any $h,h^{\prime}\in\mathcal{H}$. We note that the method introduced by [39] defines a risk functional in terms of the squared moment deviation and is thus related to GMM.
The role and interpretation of the “adversary” function $g\in\mathcal{G}$ and its associated objective function vary subtly across different saddle-point formulations, which in turn influences the types of assumptions required. This distinction is important: if $g$ is intended to approximate $\mathbb{E}\left[Y-h(X)\mathrel{}\middle|\mathrel{}Z\right]$, then the space $\mathcal{G}$ must be sufficiently rich to do so, which is reflected via a closedness condition. If $g$ is primarily a Lagrange multiplier, its existence and properties are more directly tied to conditions like a source condition. This difference helps explain why [2] can dispense with the closedness assumption while other methods may require it.
B.0.0.0.3 About the contrastive loss.
The contrastive loss used to learn spectral features derives its name from the foundational concept of contrastive learning, which traces back to the early 1990s. Notably, [5] introduced the Siamese network architecture for signature verification, which utilized a form similar to Eq.(8) to measure the distance between input pairs. Later, [9] formalized the contrastive loss function for face verification tasks. This loss is, in turn, linked to a truncated conditional operator (see Eq. (6)). The underlying principle has been frequently redeveloped under various names, such as correspondence analysis [16], principal inertia components [20, 21] for finite alphabets, the contrastive kernel [18, 11], and pointwise dependence [35] in self-supervised representation learning. The objective in Eq. (8) was also proposed and studied by [36] and used as a local approximation to the log-loss of classification deep neural networks [38]. More recently, [24] linked this same objective to the SVD of the conditional expectation operator.
Appendix C Experimental Details
In this section, we provide additional details for the experiments presented in the main text. Let $(Z,X)$ be generated as described in Section 5 with $n=100000$. Figure 3 displays both the true data-generating density and the density corresponding to a set of learned spectral features. Even for $d=11$, the resulting densities are complex. While it is feasible to conduct analogous experiments with higher values of $d$, we observed no qualitative changes in the outcomes. However, training and hyperparameter tuning became increasingly challenging as $d$ grew.
Figure 3: Comparison of a density corresponding to a set of learned spectral features (left) and the true data generating density (right).
C.1 Models Employed
The features were learned using two-hidden-layer neural networks. All models shared the same architecture, with layer widths $[1,50,50,50]$: the input is one-dimensional, and the final layer outputs $50$ learned features. To encourage the models to learn more oscillatory functions, the first layer used the activation $x\mapsto x+\sin^{2}(x)$, as introduced by [40], followed by GELU-activated hidden layers and a final linear layer.
We note that the first singular eigenfunctions of the conditional expectation operator $\mathcal{T}$ are always the constant functions (see Section 2). Therefore, we hard-code the constant feature $\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}$ into the model and restrict the following learned features to be mean-zero.
In addition, we included a regularization term to penalize both feature collinearity and large feature norms. This regularizer is a sample-based approximation of the quantity defined in Equation (10) of [24]:
$\mathbb{E}\left[\|\varphi(X)\varphi(X)^{\mathsf{\scriptscriptstyle T}}-I\|^{2}\right]+\mathbb{E}\left[\|\psi(Z)\psi(Z)^{\mathsf{\scriptscriptstyle T}}-I\|^{2}\right]+2\mathbb{E}\left[\|\varphi(X)\|^{2}\right]+2\mathbb{E}\left[\|\psi(Z)\|^{2}\right].$
The spectral features were trained on $100{,}000$ samples of $(Z,X)$, while the 2SLS estimator built from the learned features used a separate dataset of $10{,}000$ samples of $(Z,X,Y)$.
C.2 Expanded example of the ugly scenario
As noted in the main text, the difficulty of recovering $h_{0}$ increases when it is not well supported on the singular functions of the conditional expectation operator $\mathcal{T}$. If the projection of $h_{0}$ onto the span of the learned singular functions is sufficiently large, the missing components may not significantly harm the quality of the estimate. We illustrate this property with a controlled experiment designed to vary the amount of support $h_{0}$ has in the singular space of $\mathcal{T}$.
We fix $d-1$ orthonormal features $u_{i}$ for $Z$ and $v_{i}$ for $X$, and define the operator:
$\mathcal{T}=\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=1}^{d-1}\sigma_{i}u_{i}\otimes v_{i},$
mirroring the setup from Section 5. We vary the spectrum by setting $\sigma_{1:k}=c$ and $\sigma_{k+1:d-1}=0$ for $k\in\{1,\ldots,d-1\}$. We then define the target function:
$h_{0}=\frac{1}{\sqrt{d-1}}\sum_{i=1}^{d-1}v_{i},$
which is uniformly spread across the feature directions. As $k$ increases, the projection of $h_{0}$ onto the singular space of $\mathcal{T}$ increases in discrete steps from 0 to 1. This results in a corresponding qualitative improvement in the accuracy of the 2SLS estimator for $h_{0}$. Figure 4 illustrates this behavior: for small values of $k$, where $h_{0}$ is largely orthogonal to the singular functions of $\mathcal{T}$, the estimate fails to recover $h_{0}$. As $k$ grows, and more of $h_{0}$’s energy lies in the span of these singular functions, the 2SLS estimate increasingly aligns with the true function.
Figure 4: Qualitative improvement in the 2SLS estimate of $h_{0}$ as $k$ increases. When $k=1$, $h_{0}$ is orthogonal to the singular functions of $\mathcal{T}$; when $k=d-1$, it is fully contained in their span.